find the area of triangleABC if a=5, b=8, c=10
Answer by palanisamy(496)

Given, a=5, b=8 and c=10
Let s = (a+b+c)/2
s = (5+8+10)/2 = 11.5
Area of the triangle = Sqrt[s(s-a)(s-b)(s-c)]
= Sqrt[11.5*(11.5-5)(11.5-8)(11.5-10)
= Sqrt [ 11.5*6.5*3.5*1.5]
= 19.81